OpenStudy (anonymous):

Solve by factoring a) (2x -2)(x +4) = 48

OpenStudy (midhun.madhu1987):

(2x -2)(x +4) = 48 2(x - 1)(x + 4) = 48 (x - 1)(x + 4) = 24 Now open the bracket by applying distributive property...

OpenStudy (anonymous):

I do not get how to apply distributive property to this.

OpenStudy (midhun.madhu1987):